OpenStudy (anonymous):

The Endpoints of LF are L(-2,2) and F(3,1). The endpoints of JR are J(1,-1) and R(2,-3). What is the approximate difference in the lengths of the two segments?

OpenStudy (anonymous):

Could someone help with how to do not just the answer? Thx

jimthompson5910 (jim_thompson5910):

Distance from L to F (ie length of LF) d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((3-(-2))^2+(1-2)^2) d = sqrt((5)^2+(-1)^2) d = sqrt(25+1) d = sqrt(26) So the exact distance between points L and F is sqrt(26) units. Using a calculator, this approximates to 5.099 Use the formula above to calculate the distance between J and R. Once you have this distance, simply subtract.

OpenStudy (anonymous):

Thanks a bunch!